solve by addition
x-2y=-1
-x+5y=4

Answer by checkley77(12844)

x-2y=-1
-x+5y=4
-----------------
3Y=3
Y=3/3
Y=1 ANS.
X-2*1=-1
X-2=-1
X=-1+2
X=1 ANS.
PROOF:
-1+5=4
4=4